OpenStudy (anonymous):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = 16, b = 28, c = ? A. 31.46 B. 31.94 C. 32.09 D. 32.25

OpenStudy (anonymous):

@REDNECK4302 @Yuba @DollyAcquah @Destinymasha @OrangeMaster @Mr.ClayLordMath

OpenStudy (anonymous):

a^2+b^2 = C^2

OpenStudy (anonymous):

\[16^{2}+28^{2}=\]

OpenStudy (anonymous):

D

OpenStudy (anonymous):

D

OpenStudy (anonymous):

REDNECK4302 is right, but he shouldn't have just given you the answer. Use the formula I gave you next time there is a question like this.

OpenStudy (anonymous):

16*16=256 28*28=784 256+784=1040 \[\sqrt{1040}\]=32.25

OpenStudy (anonymous):

WELL AREN'T YOU A STICK IN THE MUD. BUT YOUR RIGHT NEXT TIME I WILL GIVE HER THE FORMULA TO DO IT HERSELF

OpenStudy (anonymous):

State which set of three numbers could represent the lengths of the sides of a right triangle. A. 8, 12, 16 B. 6, 7, 8 C. 16, 32, 36 D. 5, 12, 13

OpenStudy (anonymous):

You can use the formula for each set, and see which one works.

OpenStudy (anonymous):

I'll give you a hint. It's not A or B.

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

THANK YOU FOR THAT NOW I LEARNED THAT I CAN USE IT ON TRIGONMETRIC PROBLEMS BUT DO U KNOW THE FORMULA FOR THE RATE OF SPEEDOR TIME OR DISTANCE

OpenStudy (anonymous):

5*5=25 12*12=144 144+25=169 \[\sqrt{169}=13\]